### v3

# Autonomous Data Selection with Language Models for Mathematical Texts

Abstract
To improve language models’ proficiency in mathematical reasoning via continual pretraining, we introduce a novel strategy that leverages base language models for autonomous data selection. Departing from conventional supervised fine-tuning or trained classifiers with human-annotated data, our approach Autonomous Data Selection (AutoDS) utilizes meta-prompted language models as zero-shot verifiers to evaluate and select high-quality mathematical content autonomously. To demonstrate the efficacy of our method, we continuously pretrained some language models on our curated dataset, achieving substantial improvements in downstream performance on the MATH, GSM8K, and BIG-Bench Hard (BBH) tasks with a token amount reduced by orders of magnitude compared to previous continual pretraining works. Our method showcases a 2 times increase in pretraining token efficiency compared to state-of-the-art baselines, underscoring the potential of our approach in enhancing models’ mathematical reasoning capabilities. The AutoMathText dataset is available at https://huggingface.co/datasets/math-ai/AutoMathText††The code is available at https://github.com/yifanzhang-pro/AutoMathText..
1 Introduction
In the field of language modeling (Devlin et al. 2018; Radford et al. 2018, 2019; Brown et al. 2020; OpenAI 2023; Anil et al. 2023), the incorporation of domain-specific knowledge emerges as a crucial area for exploration (Lewkowycz et al. 2022; Azerbayev et al. 2023b). This is particularly important in the realm of mathematical reasoning, where the development and curation of specialized datasets for pretraining and finetuning represent a critical need and a challenge (Hendrycks et al. 2021; Paster et al. 2023; Wang, Xia, and Liu 2023). The drive toward creating language models proficient in complex mathematical reasoning underscores the importance of high-quality, domain-specific datasets. However, the mathematical field faces a scarcity of such resources, highlighting the need for innovative solutions to cultivate models with deep understanding and problem-solving skills.
Recent endeavors, such as those by Gunasekar et al. (2023) and Li et al. (2023), have made significant strides in addressing this challenge. They demonstrated the potential of leveraging GPT-4 to assess the educational value of code data within the Stack dataset (Kocetkov et al. 2022), employing model-generated annotations to train a random forest classifier for quality prediction. These studies mark a pivotal step toward enhancing the quality of data for model training. Nonetheless, they can only assign discrete labels to the data points, e.g., good or bad, instead of assigning continuous real scores, e.g., a data point of educational value vs a data point of value .
As we will demonstrate later, computing real-valued scores for training data can significantly improve the pretraining token efficiency because the model can focus on the most informative data points, where “informative” is defined by a scoring threshold. However, generating scores can be difficult for large language models (LLMs), as it has been observed that LLMs are not good at accurately generating numbers or sampling from complex distributions (Hopkins, Renda, and Carbin 2023; Hu et al. 2023). Inspired by the innovative DPO method (Rafailov et al. 2023), we propose leveraging the logits of specific tokens to directly formulate a quantitative score function, circumventing the need for extensive data labeling or classifier training.
In this work, we introduce a strategy that utilizes the intrinsic capabilities of base language models, equipped with zero-shot meta-prompts, to autonomously evaluate the mathematical quality and educational value of content. Our score function offers a more nuanced and granular analysis, unlike previous methods that primarily focused on binary classification (Li et al. 2023; Paster et al. 2023). This enables a refined and sophisticated training strategy that extends beyond the limitations of binary filtering.
The core of our contribution lies in the autonomous content evaluation without the necessity for alignment with human-labeled scores through Supervised Fine-Tuning (SFT), Reinforcement Learning from Human Feedback (RLHF) (Ouyang et al. 2022), or Direct Preference Optimization (DPO) (Rafailov et al. 2023). By employing a softmax function over logits for ‘YES’ and ‘NO’ tokens, our method autonomously assesses content relevance and value. This facilitates an active learning process where the model customizes its learning journey by querying the educational merit of materials. This approach signifies an attempt towards the realization of autonomous learning systems that are dynamic, proactive, and capable of self-directed evaluation and learning, especially in specialized fields like mathematics.
Our contributions can be listed as three-fold:
We showcase the efficacy of leveraging base language models with meta-prompts for zero-shot verification using a straightforward score function derived from logits. Our method, Autonomous Data Selection (AutoDS) advances beyond traditional alignment strategies such as SFT and RLHF without the reliance on human-annotated data, facilitating autonomous content evaluation.
We address the shortage of labeled high-quality mathematical training resources by introducing the open-source dataset on mathematical texts. This comprehensive dataset is designed to enrich AI model training with mathematical content, thereby enhancing their performance in math-intensive tasks.
Through empirical evidence, we demonstrate the effectiveness of our methodology by continuously pretrain language models on the AutoMathText dataset. Our results highlight substantial improvements in downstream performance on the MATH (Hendrycks et al. 2021), GSM8K (Cobbe et al. 2021), and BIG-Bench Hard (BBH) tasks (Suzgun et al. 2022) with 2 times pretraining token efficiency, underscoring the practical benefits of our approach in mathematical reasoning tasks.
2 Related Work
Mathematical datasets and language models.
The emergence of chain-of-thought prompting methodologies (Radford et al. 2019; Wei et al. 2022; Wang et al. 2022; Fu et al. 2022; Gao et al. 2023a; Yao et al. 2023; Zhang et al. 2023; Gou et al. 2023) has been instrumental in harnessing and enhancing the reasoning capabilities inherent within language models. Our research, however, distinctly concentrates on the domain of continual pretraining with a focus on mathematical datasets. The creation of mathematical datasets has been critical in propelling AI models’ proficiency in mathematical comprehension and reasoning. Foundational contributions, such as the AMPS dataset by Hendrycks et al. (2021) and the Proof-Pile dataset by Azerbayev et al. (2023a), have provided capstones for models to systematically tackle mathematical problems and proofs. The Llemma model (Azerbayev et al. 2023b) builds upon this foundation, dedicating its efforts to the continual pretraining of language models with mathematical data especially the OpenWebMath dataset (Paster et al. 2023), aiming to refine their complex reasoning skills further. Nevertheless, the meticulous selection of mathematical data is still an area fraught with challenges.
Data selection in language modeling.
The landscape of data selection in language modeling has seen a variety of approaches aimed at refining the quality and relevance of training data. Techniques have ranged from employing binary classifiers used by GPT-3 (Brown et al. 2020) and PaLM (Chowdhery et al. 2023) to filter web data towards more formal sources like Wikipedia and books, to more nuanced strategies that consider the difficulty or domain-specificity of the data. For example, the Minerva model (Lewkowycz et al. 2022) used rule-based filtering for mathematical content, while DSIR (Xie et al. 2023b) applied importance resampling to align the data distribution with a target domain. Furthermore, DoReMi (Xie et al. 2023a) introduces a novel angle, optimizing domain weights with a proxy model to minimize worst-case excess loss across domains. However, the low inherent perplexity (entropy) in math-related and code-related corpora suggests that DoReMi might not be optimally suited for enhancing mathematical pretraining. Recently, Gunasekar et al. (2023); Li et al. (2023) demonstrated the utility of GPT-4 in annotating data quality for the Stack dataset (Kocetkov et al. 2022), subsequently using a random forest model for classification based on these annotations. Wettig et al. (2024) propose to train a reward model called Qurating for data selecting. Our work diverges from previous approaches by introducing a fully autonomous data selection method that leverages the intrinsic capabilities of language models without the need for human-generated (and AI-generated) annotations or external trained classifiers.
Data selection across various domains.
The strategy of data selection transcends NLP tasks, extending its utility to a variety of domains, including vision and general domain adaptation. The Moore-Lewis technique, as introduced by Moore and Lewis (2010) and further refined by Axelrod (2017), exemplifies this approach by employing the cross-entropy differential between n-gram language models (LMs) tailored to specific targets and general corpora. Similarly, discrepancies in feature space and n-gram distributions have been effectively leveraged for data selection in domain adaptation scenarios, as evidenced by the work of Jiang and Zhai (2007), Liu et al. (2019), and Ruder and Plank (2017). Moreover, the significance of strategic data selection is equally acknowledged within the realm of computer vision, where methodologies aimed at optimizing training datasets have demonstrated substantial benefits. Notable contributions in this area include the pioneering curriculum learning framework by Bengio et al. (2009), the exploration of submodularity for efficient data selection by Wei, Iyer, and Bilmes (2015), and recent advancements in prioritized data selection techniques by Coleman et al. (2019) and Mindermann et al. (2022).
3 Language Models as Zero-shot Verifiers
The proliferation of language models has introduced unprecedented opportunities for advancing AI systems capable of intricate reasoning and decision-making (Wei et al. 2022; Bubeck et al. 2023). In this context, our work explores the frontier of employing base language models as zero-shot verifiers, a concept that diverges from traditional few-shot learning paradigms (Brown et al. 2020) by eliminating the need for task-specific fine-tuning or example-based prompting (Reynolds and McDonell 2021; Kojima et al. 2022; Zhang, Yuan, and Yao 2023). Our methodology embraces the zero-shot approach to leverage the inherent capabilities of language models, thereby enabling a direct assessment of textual content’s relevance and educational value in the domain of mathematics without prior alignment with human-generated labels.
Central to our approach AutoDS is the formulation of a scoring function, as delineated in Equation (1), which quantitatively evaluates the language model’s inclination towards affirming or negating the mathematical content and educational merit of a given piece of content. This function operates on the logits associated with ‘YES’ and ‘NO’ responses to meta-prompts, offering a nuanced mechanism for content evaluation:
This scoring function represents a novel integration of language models’ prediction capabilities into an autonomous evaluation framework, bypassing the limitations associated with traditional supervised learning techniques. Our approach forgoes the conventional reliance on manually labeled datasets or classifier training, instead offering a direct and nuanced assessment of content across varied mathematical sources, as exemplified in Figures 2 and 7. Figure 3 demonstrates the meta-prompt designed for autonomous data selection, illustrating how language models can evaluate the mathematical and educational value of content from diverse sources such as Common Crawl, arXiv, and GitHub (see Figure 8 and 9). Our use of meta-prompts not only serves as in-context alignment for base language models but also ensures that the language models operate within a specifically tailored syntax, enhancing their ability to produce type-safe foreseeable responses. Notice that the ‘system’ tags are directly using plain text instead of special tokens for ease of implementation without modifying the tokenizers. Responses from the model are constrained to four possibilities, thereby allowing for a refined selection process tailored to educational content in mathematics.
Leveraging the capacity for handling multiple queries within a single prompt, our methodology interprets the LM score as a pseudo-probability. This interpretation facilitates a layered assessment by aggregating the scores of individual questions. In our framework, the language model is tasked with addressing two queries simultaneously, and we derive the composite LM-Score for these inquiries utilizing Equation (2). In subsequent discussions, we refer to this aggregated measure simply as the LM-Score. This approach emphasizes the redundancy of collecting annotated data for alignment techniques like Supervised Fine-Tuning (SFT) or Reinforcement Learning from Human Feedback (RLHF), proposing a more streamlined, zero-shot in-context alignment strategy. This refined strategy not only simplifies the evaluation process but also enhances the efficiency and scalability of our AutoDS method.
Importantly, the utilization of base language models equipped with meta-prompts is instrumental in our approach, offering a highly efficient pathway for continual pretraining and active life-long learning. Through the strategic use of meta-prompts, we can tap into the innate instruction-following capabilities of these models, bypassing the need for traditional alignment mechanisms. This intrinsic property allows for the direct application of a model’s latest checkpoint to autonomously determine the suitability of data for subsequent pretraining epochs. Such a method not only streamlines the process of data curation but also ensures that the model remains dynamically attuned to the evolving landscape of mathematical content, thereby enhancing its learning trajectory and adaptability over time. This underscores the transformative potential of our approach in leveraging the existing competencies of language models for autonomous data evaluation and selection, setting a new precedent for the development of self-evolving AI systems specialized in the domain of mathematics.
Moreover, our approach deliberately avoids SFT or RLHF to anticipate and leverage the evolving superiority of language models over human evaluative capabilities, especially in domains requiring specialized knowledge like mathematics. This decision is substantiated by the examples depicted in Figures 2 and 7, which highlight the potential limitations of trained classifier-based and human-led content evaluation. OpenWebMath (Paster et al. 2023) trained a model to predict the probability a document is mathematical, which turns out not to be very satisfying (see Figure 2).
Language models, free from human biases and constraints, present a scalable and objective mechanism for content assessment, as humans may be seen as weak supervisors compared to language models themselves (Burns et al. 2023). Our methodology advocates for autonomous supervision through direct engagement by eliciting language models. This paradigm shift towards self-supervised evaluation and selection paves the way for the next generation of AI systems, characterized by their autonomous learning and adaptability in specialized knowledge domains.
4 Autonomous Data Selection with Language Models
Our study leverages three primary data sources: Common Crawl (specifically, the OpenWebMath subset (Paster et al. 2023)), arXiv (via the RedPajama dataset (Computer 2023)), and GitHub (the Stack dataset (Kocetkov et al. 2022; Azerbayev et al. 2023b)). These sources were chosen for their rich mathematical content, spanning a broad spectrum of complexity and formats.
Experiment Details. We employ the Qwen-72B base language model (Bai et al. 2023), notable for its MMLU score of 77.4, to process our datasets. Specifically, we utilize:
1.
6.32M documents from the OpenWebMath dataset (Paster et al. 2023), a curated subset of Common Crawl;
2.
1.54M documents from the arXiv subset of the RedPajama dataset (Computer 2023);
3.
3.40M documents from the Algebraic Stack dataset (Azerbayev et al. 2023b), a specialized subset of the Stack dataset.
Selection encompassing over 200GB of data, while not exhaustive, serves as a representative demonstration, prioritizing cost-effectiveness and coverage. Our computational setup includes A100-80G and A800-80G GPUs, employing the vLLM inference framework (Kwon et al. 2023) for efficient language model inference. Processing the combined 11.26M documents required approximately 750 hours on 4 A100-80G GPUs, translating to 3000 GPU hours in total. Contrastingly, manual annotation of this dataset by experts familiar with undergraduate-level and beyond mathematical content would cost upwards of $10M, assuming a rate of $1 per document. Our method significantly reduces this cost to approximately $10,000 (the cost is estimated by using Azure’s machine learning service at $3.4 per A100 GPU hour).
Visualization of Data Composition
The visualization of data composition is essential to discern the quality and diversity of the web subset of our datasets. Figure 4 displays a tree map detailing the domains by LM-Score (, ) ranges from 0.50 to 1.00 and 0.75 to 1.00, respectively. This representation not only spotlights the disparity in quality across different sources but also reveals the high-quality nature of data from StackExchange. This domain stands out, showcasing a considerable volume of content that demonstrates superior quality, yet a substantial portion of this data remains unexplored in existing literature (Wang, Xia, and Liu 2023; Liu et al. 2024), signifying a valuable opportunity for further investigation.
Delving deeper, Figure 5 offers a granular view of the LM-Score distribution across the domains. It is apparent that StackExchange, mathhelpforum.com, and physicsforums.com are leading in terms of high-quality content, with the highest proportions of scores within the 0.75 to 1.00 range. This detailed breakdown elucidates the domains where our autonomous data selection method is particularly effective, guiding the strategic direction for subsequent data preprocessing and model training efforts.
| LM-Score | Type | # Tokens (M) | MATH Acc. (CPT) (%) | MATH Acc. (SFT) (%) |
| - | Baseline (w/o pretraining) | 0 | 12.88 | 27.20 |
| - | OpenWebMath | 328.9 | 10.50 | 26.98 |
| 0.75-1.00 | AutoDS | 328.9 | 13.68 | 28.06 |
| Model & Selection Method | MATH (5-shot) | GSM8K (5-shot) | BIG-Bench Hard (3-shot) |
| Gemma-2B Base | 10.96 | 17.29 | 34.19 |
| + Uniform (OpenWebMath) | 10.16 | 18.88 | 36.34 |
| + DSIR | 5.62 | 11.90 | 34.43 |
| + QuRating | 9.76 | 13.19 | 31.76 |
| + AutoDS | 11.02 | 18.88 | 34.88 |
| LLaMA2-7B Base | 2.94 | 12.51 | 39.89 |
| + Uniform (OpenWebMath) | 5.14 | 19.79 | 41.53 |
| + DSIR | 2.56 | 12.51 | 39.49 |
| + QuRating | 2.90 | 10.54 | 39.27 |
| + AutoDS | 7.74 | 21.99 | 42.76 |
| Mistral-7B Base | 12.88 | 38.82 | 55.92 |
| + Uniform (OpenWebMath) | 14.26 | 44.12 | 56.50 |
| + DSIR | 12.30 | 42.00 | 55.97 |
| + QuRating | 12.90 | 36.32 | 55.63 |
| + AutoDS | 16.14 | 45.41 | 58.61 |
5 Experiments
In this section, we want to test the effectiveness of the AutoDS method in enhancing the mathematical reasoning capabilities of language models. To this end, we continually pretrained language models like Gemma (Team et al. 2024), LLaMA2 (Touvron et al. 2023), and Mistral (Jiang et al. 2023) showcasing the efficiency of our data selection method. Contrasting with the extensive 200B-token training performed by Llemma (Azerbayev et al. 2023b), we utilized merely less than 1.5% of that amount (less than 3B tokens), thereby emphasizing the potential of our data-efficient training approach. Our experiments include baselines employing uniform sampling, DSIR (Xie et al. 2023b), Qurating (Wettig et al. 2024), and our AutoDS method leveraging LM-Score-based selection. Token counts were balanced among different types of training data to ensure comparability.
Experiments on Continual Pretraining
Experiment details. Utilizing the LLaMA-Factory (hiyouga 2023), we perform the continual pretraining of the Mistral-7B-v0.1 model with three epochs, using a cosine learning rate schedule with a 3% warm-up period and a peak learning rate of 5e-6. The DeepSpeed framework (Rajbhandari et al. 2020) with ZeRO-2 Stage optimization facilitates our training acceleration. The models are continual-pretrained on a node comprising 8xA800 GPUs. We set the total batch size to 256. We first utilize the selected data from the web subset with the highest quality for a preliminary evaluation.
Evaluation results. Our evaluation protocol adheres to the standard eval harness framework (Gao et al. 2023b), consistent with the Huggingface Leaderboard’s protocol (Beeching et al. 2023). The results, as detailed in the tables below, illuminate the efficacy of our AutoDS dataset in enhancing the model’s performance.
| NQ | ARC-E | ARC-C | SciQ | LogiQA | BoolQ | |||
| Model & Selection Method | (5) | (5) | (25) | (25) | (2) | (2) | (0) | Average |
| Gemma-2B Base | 14.88 | 36.60 | 77.61 | 46.50 | 96.30 | 25.35 | 69.54 | 42.92 |
| + Uniform (OpenWebMath) | 13.80 | 36.54 | 77.40 | 45.39 | 96.40 | 26.27 | 68.35 | 42.95 |
| + DSIR | 13.27 | 34.44 | 77.27 | 45.82 | 95.90 | 23.50 | 54.92 | 39.71 |
| + QuRating | 14.32 | 33.81 | 77.95 | 46.76 | 96.20 | 24.42 | 68.53 | 41.67 |
| + AutoDS | 13.27 | 36.09 | 76.81 | 46.08 | 96.10 | 27.19 | 71.28 | 43.16 |
| LLaMA2-7B Base | 26.01 | 37.08 | 80.72 | 49.74 | 96.80 | 26.57 | 77.68 | 44.99 |
| + Uniform (OpenWebMath) | 26.07 | 40.09 | 80.77 | 50.09 | 96.70 | 27.65 | 78.41 | 46.62 |
| + DSIR | 25.76 | 36.63 | 80.98 | 48.98 | 96.50 | 26.73 | 72.54 | 44.27 |
| + QuRating | 25.96 | 37.84 | 80.43 | 50.60 | 96.80 | 27.65 | 77.71 | 44.97 |
| + AutoDS | 25.84 | 40.66 | 80.09 | 49.74 | 96.70 | 27.96 | 77.19 | 47.07 |
| Mistral-7B Base | 29.81 | 52.39 | 84.68 | 57.25 | 97.40 | 30.26 | 83.58 | 54.30 |
| + Uniform (OpenWebMath) | 29.17 | 52.17 | 84.18 | 56.66 | 97.20 | 31.03 | 83.82 | 54.91 |
| + DSIR | 29.22 | 52.62 | 84.72 | 57.25 | 97.30 | 30.26 | 73.76 | 53.54 |
| + QuRating | 28.89 | 52.01 | 85.48 | 57.76 | 97.30 | 31.18 | 82.81 | 54.03 |
| + AutoDS | 29.06 | 52.30 | 84.18 | 55.20 | 96.80 | 31.03 | 83.12 | 55.19 |
In Table 1, we compare the MATH test accuracy of models after continual pretraining. The auto-selected data consistently outperforms its uniform counterpart, achieving higher accuracy percentages. Notice that the uniformly sampled data from the OpenWebMath dataset have already been filtered using OpenWebMath’s rule-based filter and trained classifier. This enhancement in performance highlights the strategic advantage of using high-quality, domain-specific data for continual model pretraining. Table 1 further examines the MATH test accuracy after supervised fine-tuning (SFT) on the MetaMathQA (Yu et al. 2023) dataset. In this SFT setting, the auto-selected data models again exhibit superior accuracy, affirming the robustness of our pretraining approach. These results underscore the AutoDS dataset’s ability to enhance model performance and as a foundation for subsequent fine-tuning processes.
Comparison with Baselines
In this subsection, we conduct experiments at a larger scale to comprehensively evaluate different data selection methods. Specifically, we continually pretrained Gemma-2B, LLaMA2-7B, and Mistral-7B models for one epoch with approximately 2.5 billion tokens.
Experiment details. In this experiment, we use a constant learning rate of 1e-6 for Mistral-7B and 1e-5 for others as the default for all methods for fair comparisons. For the OpenWebMath dataset (Paster et al. 2023), we use uniform sampling as a baseline, notice that the OpenWebMath has already been selected using trained classifiers. The AutoDS method’s data selection ranged from LM-Scores of 0.6 to 1.0 within the web subset. For the DSIR method, it requires a target dataset to calculate the KL divergence between the source dataset and the target dataset. We use the Pile (Gao et al. 2020) dataset’s wiki validation set as the target dataset. For the Qurating method (Wettig et al. 2024), we directly utilized the QuratedPajama dataset, selecting data based on top-k scores of educational value.
Experimental results. In Figure 1, the checkpoint evaluation every 100 steps, approximately 52 million tokens. From Figure 1 and Table 2, models pretrained with the data selected using the AutoDS method consistently show superior performance across a diverse set of complex reasoning tasks including MATH (Hendrycks et al. 2021), GSM8K (Cobbe et al. 2021), and BIG-Bench Hard (Suzgun et al. 2022), highlighting the method’s robustness and the AutoDS dataset’s effectiveness in enhancing models’ reasoning capabilities. Notably, on the MATH dataset, AutoDS shows 2.36 times pretraining token efficiency compared to the OpenWebMath uniform sampling baseline (14.26%), achieving 16.14% accuracy with only 2.5B tokens for continual pretraining.
Beyond complex reasoning, we extended our evaluation to assess how well the models adapted to other cognitive domains, such as commonsense reasoning, world knowledge, and reading comprehension. Table 3 encapsulates this multi-faceted performance evaluation. It’s noteworthy that while AutoDS did not universally top all categories, its average performance across diverse tasks (including all tasks shown in Table 2 and Table 3) shows the superiority of our method compared to other data selection methods. These outcomes strongly advocate for the AutoDS’s potential to advance language models in mathematical reasoning and beyond.
6 Conclusion
Our method leverages the inherent self-evaluation and active learning capabilities of language models significantly improving the quality and relevance of training data in intricate and specialized fields like mathematics. This research opens the door to further investigations into autonomous data curation and model training techniques, heralding a new era in AI’s capacity for understanding, reasoning, and innovation within specialized domains.
Ethical Statement
This study, aimed at enhancing the capabilities of language models through autonomous data selection and continual pretraining, presents insightful implications for the field of AI research, particularly in the training and development of language models with specialized knowledge. The deployment of autonomous systems for the selection of training data introduces considerations of transparency, fairness, and accountability within the AI development process. By reducing reliance on human-labeled data, our method shifts the responsibility for content evaluation to the AI itself, raising important questions about the model’s decision-making processes. Ensuring these processes are transparent and free from biases is essential to prevent the perpetuation of existing inequalities or the introduction of new biases into AI systems.
Appendix A Appendix for Examples
Web Subset
Code Subset
Arxiv Subset
Appendix B More on Experiments
More Experimental Results
| Model & Selection Method | HellaSwag (10-shot) | PIQA (6-shot) | WinoGrande (15-shot) |
| Gemma-2B Base | 48.30 | 70.67 | 60.54 |
| + Uniform (OpenWebMath) | 52.91 | 76.71 | 66.38 |
| + DSIR | 52.95 | 77.15 | 66.61 |
| + QuRating | 53.10 | 77.53 | 66.38 |
| + AutoDS | 52.82 | 77.42 | 66.61 |
| LLaMA2-7B Base | 58.88 | 79.43 | 75.85 |
| + Uniform (OpenWebMath) | 58.43 | 79.54 | 75.30 |
| + DSIR | 58.38 | 78.84 | 75.37 |
| + QuRating | 58.79 | 79.00 | 74.66 |
| + AutoDS | 58.28 | 78.18 | 74.51 |
| Mistral-7B Base | 62.82 | 82.10 | 81.22 |
| + Uniform (OpenWebMath) | 62.21 | 82.21 | 80.19 |
| + DSIR | 63.10 | 81.94 | 81.37 |
| + QuRating | 62.64 | 81.99 | 80.11 |
| + AutoDS | 62.72 | 82.21 | 80.03 |
Prompts
Alternative Score functions
One can use alternative scoring functions corresponding to different partition functions, such as the formulas shown below.
Or: